OpenStudy (azureilai):

The initial dimensions of a rectangle are 5x7 cm and the length and width are increasing at 2cm/second. Find an algebraic functions to represent area, A, of the rectangle as a function of the time, t.

OpenStudy (gorv):

length= 5+2t width=7+2t area=(5+2t)*(7+2t)=35+10t+14t+4t^2=4t^2+24t+35

OpenStudy (gorv):

@Azureilai

OpenStudy (azureilai):

I see, thank you. I was confused on how to incorporate time into there.